OpenStudy (anonymous):

Which choice is a factor of x^4-1 when it is factored completely? a.) x^2-1 b.) x-1 c.)x d.)1

zepdrix (zepdrix):

Remember how to factor the difference of SQUARES? c: \[\large a^2-b^2=(a-b)(a+b)\] We can manipulate this thing to make it look like the difference of squares.

OpenStudy (anonymous):

^-^ mmhmm

zepdrix (zepdrix):

\[\large x^4-1\qquad\rightarrow\qquad x^4-1^4 \qquad\rightarrow\qquad (x^2)^2-(1^2)^2\]Now we have the difference of squares. Applying the rule gives us,\[\large \left(x^2-1^2\right)\left(x^2+1^2\right)\]

zepdrix (zepdrix):

Now the problem has some tricky little wording :) You might look at your factors and go "AHA! I see \(x^2-1\)!" But the instructions ask you for a factor after you've COMPLETELY factored it. We can actually factor this a bit further.

zepdrix (zepdrix):

because \(x^2-1^2\) is also the difference of squares! Let's apply the rule again.

zepdrix (zepdrix):

\[\large \color{orangered}{\left(x^2-1^2\right)}\left(x^2+1^2\right)\qquad\rightarrow\qquad \color{orangered}{(x-1)(x+1)}\left(x^2+1\right)\]

zepdrix (zepdrix):

NOW it's finally fully factored :D Do any of those factors match your choices? If you're confused by any step, let me know :U

OpenStudy (anonymous):

B?

zepdrix (zepdrix):

ya

OpenStudy (anonymous):

So its B :o